OpenStudy (anonymous):

Hi everyone simple question i really need help with What is the exact value of the square root of 28 over 9 the square root of 16.? Simplify if possible. Answer the square root of 7 over 9. 2 over 9. the square root of 7 over 18. 2 over 18.

OpenStudy (lgbasallote):

can you draw it?

OpenStudy (anonymous):

like the equation

OpenStudy (helder_edwin):

is this correct? \[ \LARGE \frac{\sqrt{28}}{9\sqrt{16}} \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the answers if you want

OpenStudy (helder_edwin):

then \[ \LARGE \frac{\sqrt{28}}{9\sqrt{16}}=\frac{\sqrt{4\cdot 7}}{9\sqrt{4\cdot 4}} \] can you go on?

OpenStudy (anonymous):

sqrt(28) ----- 9 sqrt(16)

OpenStudy (helder_edwin):

no! apply the properties of roots (distribution of powers to products)

OpenStudy (anonymous):

hmm

OpenStudy (helder_edwin):

seriously! \[ (ab)^n=a^nb^n \] so \[\large \frac{\sqrt{4\cdot7}}{9\sqrt{4\cdot4}}=\frac{\sqrt{4}\sqrt{7}}{9\sqrt{4}\sqrt{4}} =\frac{\sqrt{7}}{9\sqrt{4}} \]